OpenStudy (anonymous):

evaluate 2x^-1y^3 if x=2 and y=-1

OpenStudy (anonymous):

insert 2 instead of x and insert -1 instead of y

OpenStudy (anonymous):

\[2(2)^{-1}-1^{3}\]

OpenStudy (anonymous):

like that now what do i do

OpenStudy (anonymous):

\[2(2)^{-1}(-1)^{3}\]

OpenStudy (anonymous):

\[n^{-1}=\frac{1}{n}\] \[n^3=n\cdot n\cdot n\]

OpenStudy (anonymous):

okay i understand that part i just need help with the combined part

OpenStudy (anonymous):

so just multiply and subtract answers you get

OpenStudy (anonymous):

(\[1/4* 1^{-3}\]

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

\[x^{-1}= 1/x\] so 

OpenStudy (anonymous):

wait \[1/4(1^{3})\]

OpenStudy (anonymous):

plugging your numbers

OpenStudy (anonymous):